OpenStudy (anonymous):

Idk I think i may just ::scream::; I have another annoying problem that I simply cant figure out

OpenStudy (anonymous):

Find the vector flux of G=2yi+3xj-zk through the cylinder x^2+y^2=4 and \[2\le z \le -2\]

OpenStudy (amistre64):

the cylindar is again simple enough for us; the normal is <x/2,y/2,0>

OpenStudy (amistre64):

the rotation is again from 0 to 2pi

OpenStudy (amistre64):

\[\int_{0}^{2pi}\int_{-2}^{2}G.N\ adzdt\]

OpenStudy (amistre64):

where a is our Normal length .... for some reason

OpenStudy (amistre64):

G= <2y, 3x, -z> N=<x/2, y/2, 0> ---------------- xy+3xy/2+0 = 5xy/2 5xy/2 * 2 = 5xy x=2cos(t), y=2sin(t) \[\int\int20sin(t)cos(t)dzdt\]

OpenStudy (amistre64):

and with any luck, i did that correctly :)

OpenStudy (amistre64):

20 sin cos = 10 sin(2t) if it helps

OpenStudy (amistre64):

\[5\int\int 2sin(2t)dzdt\]

OpenStudy (amistre64):

we agree?

OpenStudy (anonymous):

kkk got to that point

OpenStudy (anonymous):

yes i agree LOL Now i am finally getting it YAY

OpenStudy (amistre64):

im just remembering it from that video i linked you to last time :)

OpenStudy (anonymous):

hmmm like y did u keep the 2 in? like i know its an identity but like its a constant so we can remove it

OpenStudy (amistre64):

yeah, i was thinking it be easier in to integrate sin2t with; but if we do z first its kinda pointless to begin with

OpenStudy (amistre64):

4sin2t +4sin2t = 8sin2t pull out the 4 and leave the 2 20 int 2sin2t dt = 20(-cos(2t)), from 0 to 2pi

OpenStudy (amistre64):

-20 cos(4pi) - - 20cos(0) = 0 .... we got a answer key for this one? to dbl chk

OpenStudy (anonymous):

noppppppeeee not this time. Its not from the textbook :(

OpenStudy (amistre64):

hmm, id have to review that video one last time to see that im remembering it correctly

OpenStudy (anonymous):

it seems to me to be correct

OpenStudy (anonymous):

Like we followed teh same process as the previous example

OpenStudy (amistre64):

ok, i reviewed the tape ... they still call it a tape right?

OpenStudy (anonymous):

lol ummmm tape=cassette

OpenStudy (amistre64):

G= <2y, 3x, -z> N=<x/2, y/2, 0> ---------------- xy+3xy/2+0 = 5xy/2 this is good still the x=2c and y=2s is good as well 5.2c.2s ------- = 10 s c = 5 s 2t 2

OpenStudy (anonymous):

hmmm i didnt follow this to be honest

OpenStudy (anonymous):

ohh its 5sin(2t) k got that

OpenStudy (amistre64):

\[\int_{0}^{2pi}\int_{-2}^{2}5sin(2t)\ 2\ dzdt\] \[\int_{0}^{2pi}\int_{-2}^{2}10sin(2t)\ \ dzdt\] \[\int_{0}^{2pi}10(2)sin(2t)-10(-2)sin(2t)\ \ dt\] \[\int_{0}^{2pi}40sin(2t)\ \ dt\] \[-20cos(4pi)+20cos(0)=0\]

OpenStudy (amistre64):

i keep getting zero :)

OpenStudy (anonymous):

Then 0 it is:D

OpenStudy (anonymous):

Thanks amistre like u saved the day

OpenStudy (amistre64):

lol, only if its right ;)